An electron and a proton are separated by a distance of 1.8 x1010 m (roughly the diameter of a single atom). What is the magnitude of the electric interaction between the electron and the proton? Number magnitude of electric foroeL What is the magnitude of the gravitational interaction between the electron and the proton? Number magnitude of gravitational force- n this scanario, how many limes stonger is the lectic orce thanh gravdational ore? Number The electric force is times stronger than the gravitational force.
Answer #1

a)

Fe = k q1q2 / r^2

Fe = 9*10^9* (1.6*10^-19)^2 / ( 1.8*10^-10)^2*

Fe = 7.11*10^-9 N

=======

b)

Fg = G m1 m2 / r^2

Fg = 6.67*10^-11* 1.67*10^-27*9.1*10^-31 / ( 1.8*10^-10)^2

Fg = 3.1285*10^-48 N

=======

c)

Fe / Fg = 7.11*10^-9/ ( 3.1285*10^-48)

Fe / Fg = 2.27*10^39

========

do comment in case any doubt, will reply for sure.. Goodluck